OpenStudy (anonymous):

Astronomers measure large distances in light-years. One light-year is the distance that light can travel in one year, or approximately 5.88 × 1012 miles. Suppose a star is 9.8 × 101 light-years from Earth. In scientific notation, approximately how many miles is it?

OpenStudy (anonymous):

5889864.288

OpenStudy (anonymous):

in scientic notation its 5.889864288x10^6

OpenStudy (anonymous):

I messed up its really suppose to be 5.88 x 10^12 and 9.8 x 10^1

OpenStudy (anonymous):

and it's either these answers 5.88 × 1013 miles 5.76 × 1014 miles 5.88 × 1012 miles 9.8 × 1012 miles

OpenStudy (xishem):

\[9.8*10^{1}lightyears*\frac{5.88*10^{12}mi}{1lightyear}=5.76*10^{14}mi\]